OpenStudy (anonymous):

Michael took his younger brother to the movies. Michael paid regular price, but it cost $2 less for his brother at the child price. It cost Michael $12 total for the movie. How much did it cost for his brother?

OpenStudy (anonymous):

Micheal payed 12$ total for the movie. Meaning he spent 12$ on his ticket and his brother's ticket. M= Price of Micheal's Ticket B = Price of brother's ticket. [M+B=12] We know that his brother's ticket is 2$ less so [M-2 = B]. Input equation 2 into equation 1.[M+(M-2)=12]. Solve M=7. Then solve for B. Micheal's brother's ticket should cost $5.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

No problem !